6) Graph by using the first and second derivative tests of f(x) = x +3x2 - 9x +4
Answer #1

If any confusion Please comment, I will help you, Thank you

DATE: fx)= x3 + 3x² - 9x + 4 I(x) = 3x² + 6x-9 3x2+6x-9-0 X -3, 1 X=-3 fw = 6X76 f^(-3) = EX(-3). + 6 = -18+6 = So, maximumDATE: x² + 2x - 320 - 3cxc1 YC (-3D бар» 1(0.57,0) 2(1.395,0) (4.97, 0)